OpenStudy (anonymous):

F(x)= 2x^2-2x, G(x)= 3x+3 H(x)= x-1 1.) f(x)-g(x) 2.) h(x) * g(x) 3.) f(x) / g(x) 4.) f(x) / h(x) simplified 5.) h(g(x)) 6.) g(f(x)) show work

OpenStudy (anonymous):

??

OpenStudy (zepp):

Basically, you take what function is represent, and do the calculation. For example; 1.) \(f(x)-g(x)\) We have \(f(x)= 2x^2-2x\), \(g(x)= 3x+3\) Then \(f(x)-g(x)=(2x^2-2x)-(3x+3)\)

OpenStudy (anonymous):

anwsers?

OpenStudy (zepp):

Do it yourself!

OpenStudy (anonymous):

:P

OpenStudy (zepp):

3.) \(\large \frac{f(x)}{g(x)}\) For this one you have to factorize \(f(x)\)

OpenStudy (anonymous):

arrrrrgh

OpenStudy (zepp):

For this one, \(h(g(x))\), usually, stuffs between parentheses is what you are going to replace x, in this case, replace x with \(g(x)\).

OpenStudy (zepp):

\(h(x)= x-1\); \( g(x)= 3x+3\). \(f(g(x))=(3x+3)-1=3x+2\)

OpenStudy (zepp):

Good luck! :D

OpenStudy (anonymous):

1. (2x^2-2x)-(3x+3) 2. (x-1)(3x+3) 3. (2x^2-2x)/(3x+3) 4. (2x^2-2x)/(x-1) 5. (3x+3)-1 6. 3(2x^2-2x)+3 Simplify these from there.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

1.) -2x- 6x = -8?

OpenStudy (anonymous):

There's no factoring in 1. All you need to to is subtract. 2x^2-2x-3x-3 (distributing the negative) 2x^2-5x-3

OpenStudy (anonymous):

what DOES that mean?

OpenStudy (anonymous):

What does what mean?

OpenStudy (anonymous):

im stupid

OpenStudy (anonymous):

Do you get how to do 1?

OpenStudy (anonymous):

forget it

OpenStudy (precal):

we will help you, you do need to respond back when asked to .